### v1

# Comment on the ”Electric Power Generation from Earth’s Rotation through its Own Magnetic Field”

Abstract
The suggestion made by C. F. Chyba and K. P. Hand about electric power generation from Earth’s rotation through its own magnetic field is intriguing [1, 2]. Here, we reanalyze the theoretical arguments for their conclusion, especially those presented in [1]. The model that they consider, is a magnetic cylindrical shell moving with low velocity $v$ in the $y$ direction at a right angle to the direction of the Earth’s magnetic field $B_{\infty}$. We first focus on one of the electromagnetic boundary conditions for moving dielectric boundaries (a topic not dealt with in [1, 2]), and obtain therefrom a result at variance with the main result in the mentioned papers. Therefore, we have to conclude that their result does not satisfy this boundary condition. Moreover, by a simple calculation, we observe that a displacement of the shell in the mentioned $y$ direction, does not involve any expenditure of external work. A process involving dissipation of energy as envisaged in [1, 2] thus appears to be forbidden for energy conservation reasons.
I Introduction
[…]
There have appeared some criticisms of this idea - cf. for instance the paper of Jeener [4] and its rebuttal in [5]. It is also notable that in an earlier experiment of Veltkamp and Wijngaarden no energy production of this magnitude was found [6]. As expressed by Wijngaarden in [3]: ”I am still convinced that the theory of Chyba et al. cannot be correct”.
In view of the uncertainty of the situation, we decided to make a reconsideration of the theory presented in [1] and [2], although their theory is detailed and very carefully done. Mainly, we will focus on their first paper [1]. Actually, we find that there is a missing piece in the formalism, namely the electromagnetic boundary condition on dielectric surfaces in motion. We discuss the implications of this omission for the main conclusion of the whole analysis, as our results are at variance with each other. We also find that the principle of energy conservation gives an important restriction in this case.
The experimental results reported in [2] will not be further considered here. To attempt a closer scrutiny of this would be complicated and lead us far away from the simple intention of the present note.
We start in the next section with the static case $\bf v=0$, and consider thereafter the case where $\bf v\neq 0$.
II The case $\bf v=0$
The geometry of the setup is clearly sketched in [1] and [2], and will not be reproduced here. It will however be convenient to reproduce the expressions for the components of the static magnetic field, called ${\bf B}_{0}$, both outside and inside the dielectric shell, in terms of cylindrical coordinates $\{\rho,\phi,z\}$,
$B_{0x}(\rho>b)=B_{\infty}+\beta_{3}(b/\rho)^{2}\cos 2\phi,$ (1)
$B_{0y}(\rho>b)=\beta_{3}(b/\rho)^{2}\sin 2\phi,$ (2)
$B_{0x}(a\leq\rho\leq b)=\beta_{1}-\beta_{2}(a/\rho)^{2}\cos 2\phi,$ (3)
$B_{0y}(a\leq\rho\leq b)=-\beta_{2}(a/\rho)^{2}\sin 2\phi,$ (4)
$B_{0x}(\rho\leq a)=2\beta_{1}(\mu_{r}+1)^{-1},$ (5)
$B_{0y}(\rho\leq a)=0.$ (6)
Here $y=\rho\sin\phi$, and the expressions for the coefficients are
$\beta_{1}=2B_{\infty}\mu_{r}(\mu_{r}+1)\zeta,$ (7)
$\beta_{2}=2B_{\infty}\mu_{r}(\mu_{r}-1)\zeta,$ (8)
$\beta_{3}=B_{\infty}[1-(a/b)^{2}](\mu_{r}^{2}-1)\zeta,$ (9)
$\zeta=[(\mu_{r}+1)^{2}-(a/b)^{2}(\mu_{r}-1)^{2}]^{-1}.$ (10)
The magnetic vector potential is directed along the $z$ axis, $^{\bf A_{0}}=A_{0}\hat{\bf z}$, where
$A_{0}(\rho>b)=B_{\infty}y+\beta_{3}(b^{2}/\rho)\sin\phi,$ (11)
$A_{0}(a\leq\rho\leq b)=\beta_{1}y-\beta_{2}(a^{2}/\rho)\sin\phi,$ (12)
$A_{0}(\rho<a)=2\beta_{1}(\mu_{r}+1)^{-1}\rho\sin\phi.$ (13)
III The case $\bf v\neq 0$}
When the shell moves with constant velocity $\bf v$ in the $y$ direction, the Maxwell equations within the shell will be modified. We call the moving frame of reference $K$, while the static reference system considered in the previous system is called $K_{0}$. Since the moving case involves energy dissipation, the conductivity $\sigma$ of the material will have to be accounted for. In analogy to the well known Reynolds number in hydrodynamics, in Refs. [1, 2] was defined a magnetic Reynolds number, $R_{m}=v\xi/(\sigma\mu)^{-1}$, where $\xi$ is a typical length over which the current flow changes appreciably, and $(\sigma\mu)^{-1}$ is the analog of the hydrodynamic kinematic viscosity. It is natural to associate $\xi$ with the outer radius $b$.
A central quantity in this kind of theory is the vector potential $\bf A$. It is pointing in the $z$ direction, $^{\bf A}=A_{z}\hat{\bf z}$, in the frame $K$ as well as in $K_{0}$. The governing equation for $A_{z}$ in $K$ is (cf. Eq. (25) in [1]),
$\partial A_{z}/\partial t+v\partial A_{z}/\partial y=\eta\nabla^{2}A_{z},$ (14)
where $\eta=(\sigma\mu)^{-1}$ is conventionally called magnetic diffusivity. The equation above is solved in two steps,
$A_{z}=A_{z}(\rho,\phi)+A_{t}(\rho,\phi,t),$ (15)
where $A_{z}(\rho,\phi)$ relates to the steady state equation
$v\partial A_{z}/\partial y=\eta\nabla^{2}A_{z},$ (16)
and $A_{t}(\rho,\phi,t)$ relates to the time dependent equation
$\partial A_{t}/\partial t+v\partial A_{t}/\partial y=\eta\nabla^{2}A_{t}.$ (17)
The solution found in [1, 2] for the latter equation decays exponentially in time,
$A_{t}(\rho,\phi,t)=C_{0}\,m(\rho)n(\phi)\exp{(k\rho\sin\phi)}\cdot\exp{[-\eta(k^{2}+\lambda^{2})t]}.$ (18)
Here $C_{0}$ is a constant, $m(\rho)$ and $n(\phi)$ are linear combinations of Bessel functions and trigonometric functions respectively, and $k$ is defined as $k=v/2\eta$, thus a $K$-dependent quantity. The constant $\lambda^{2}$ is related to a separation constant in Eq. (17). As emphasized in [1, 2], the time-dependent part of the solution decays swiftly with the assumed low value of the Reynolds number $R_{m}$, and can be ignored when seeking for stationary solutions.
The time-independent solution given, (Eq. (49) in [1]), is
$A_{z}(a\leq\rho\leq b)=k\beta_{1}z^{2}+\beta_{1}y-\beta_{2}ka^{2}K_{1}(k\rho)e^{ky}\sin\phi,$ (19)
where $K_{1}$ is a Bessel function. From this the gradient of the scalar potential is derived as $\nabla V=-v\beta_{1}\hat{\bf z}$. Correspondingly, the axial electric field in the interior of the shell, i.e. ${\bf E}=-\nabla V-\partial{\bf A}/\partial t$ becomes
${\bf E}(a\leq\rho\leq b)=v\beta_{1}\hat{\bf z}$, (20)
when omitting the last term. The electric field on the outside of the shell is obtained by setting $\mu_{r}=1$, whereby $\beta_{1}=B_{\infty}$ and so
${\bf E}(\rho>b)=vB_{\infty}\hat{\bf z}$. (21)
There are two boundary conditions for dielectric boundaries moving with small velocities [7]. When applied to $\rho=b$ they read
${\bf n}\times[{\bf E}(b+)-{\bf E}(b-)]=v_{n}[{\bf B}(b+)-{\bf B}(b-)],$ (22)
${\bf n}\times[{\bf H}(b+)-{\bf H}(b-)]=-v_{n}[{\bf D}(b+)-{\bf D}(b-)],$ (23)
where $v_{n}$ is the velocity of the boundary along the normal $\bf n$. We shall focus only on Eq. (22). Inserting the expression above for $\bf E$ on the left hand side, we first get
${\bf n\times\hat{\bf z}}]v(B_{\infty}-\beta_{1})=v_{n}[{\bf B}(b+)-{\bf B}(b-)].$ (24)
The fields ${\bf B}(b+)$ and ${\bf B}(b-)$ are thus tangentially directed. Taking the $x$ components, we obtain
$B_{\infty}-\beta_{1}=B_{x}(b+)-B_{x}(b-);$ (25)
the two $v$ factors on each side drop out.
We want to solve this equation with respect to the component $B_{x}(b-)$ on the inside surface. For that we need the expression for the exterior component $B_{x}(b+)$ also. This component was not given explicitly in [1, 2], but can be derived as follows, making use of the same method. We start from Eq. (11), modifying the last term, $A_{0}(\rho>b)=B_{\infty}y+h(z)$, where $h(z)$ is determined from Eq. (16) to be $h^{\prime\prime}(z)=2kB_{\infty},$ or $h(z)=kB_{\infty}z^{2}.$ Going to the case $v\neq 0$, we get the extended form
$A_{z}(\rho>b)=kB_{\infty}z^{2}+B_{\infty}y+C\cdot K_{1}(k\rho)e^{ky},$ (26)
with $C$ a constant, and $K_{1}(k\rho)=1/(k\rho)$ approximately for small values of $k\rho$. The value of $C$ is determined from requiring agreement with the expression (11) for $v=0$ for small values of $ky$, implying $C=\beta_{3}kb^{2}$, leading to the external potential
$A_{z}(\rho>b)=kB_{\infty}z^{2}+B_{\infty}y+\frac{\beta_{3}b^{2}}{\rho}\sin\phi(1+k\rho\sin\phi).$ (27)
Comparing this with the $v=0$ expression for $A_{0}$, we can write it in the form ($z^{2}$ term omitted)
$A_{z}(\rho>b)=A_{0}(\rho>b)+\frac{1}{2}\beta_{3}R_{m}b\sin\phi,$ (28)
where
$R_{m}=\mu\sigma vb$ (29)
is the magnetic Reynolds number. In a vacuum, $\sigma=0$, so we can just identify $A_{z}$ with $A_{0}$ on the outside.
Actually, this could have been seen in a more direct way, by exploiting the gauge condition $\nabla\cdot A=0$, which was found to apply also when $v\neq 0$ (cf. Eq. (22) in [1] in the vacuum case $\eta\rightarrow\infty$.
Then we can make use of the expression (1) directly in the frame $K$, implying that
$B_{0x}(b+)=B_{\infty}+\beta_{3}\cos 2\phi.$ (30)
Thus the final result, in the same frame $K$, becomes
$B_{x}(b-)=B_{0x}(b+)+\beta_{1}-B_{\infty}=\beta_{1}+\beta_{3}\cos 2\phi.$ (31)
This expression is independent of $v$. It can be compared with the expression presented in [1] (Eqs. (59) and (60)),
$B_{x}(b-)|_{CH}=B_{0x}(b+)-R_{m}\beta_{2}(a/b)^{2}\sin\phi\cos^{2}\phi,$ (32)
where, for low $v$,
$R_{m}=2kb=\mu\sigma vb.$ (33)
Thus, our application of the electrodynamic boundary conditions for moving boundaries leading to the expression (31), does not agree with the expression (32) from Ref. [1]. Although this fact is disturbing, it does not in opinion solve the essence of the problem. The main point is rather that a process as envisaged herein cannot occur at all, due to energy conservation. This will be elaborated upon in the next section.
Therefore, our inclusion of the electromagnetic boundary conditions for moving boundaries leading to the expression (31), disagrees with the expression (32) given in Ref. [1].
IV Discussion and Summary
We go back to the static case $v=0$, and calculate the total magnetic force component $F_{y}$ on the shell, per unit length in the $z$ direction. To this end we only need the Maxwell stress tensor $T_{ik}$ on the outside, at $\rho=b+$. An important point is that the surface stress originates from the volume force density ${\bf f}=-(1/2)H^{2}\nabla\mu$ acting in the interior of the dielectric boundary layer. Upon integrating the normal component of this force across the boundary region one obtains the surface pressure in its conventional form, as a difference between the normal components of the Maxwell tensor on the two sides of the surface. This force always acts towards the optically thinner region, which in our case with $\mu_{r}>1$ means that it acts outwards (cf., for instance, [8, 9, 10]).
For $\rho>b$, the only component of the Maxwell tensor needed is
$T_{\rho\rho}=\frac{1}{\mu_{0}}B_{0\rho}^{2}-\frac{1}{2\mu_{0}}{\bf B}_{0}^{2}.$ (34)
From Eqs. (1) and (2) we have on the outer surface
$B_{0\rho}(\rho+)=(B_{\infty}+\beta_{3})\cos\phi,$ (35)
${\bf B}_{0}^{2}(\rho+)=B_{\infty}^{2}+2B_{\infty}\beta_{3}\cos 2\phi+\beta_{3}^{2}.$ (36)
Inserting these expressions into the integral we obtain
$F_{y}=\int_{0}^{2\pi}T_{\rho\rho}(b+)\rho\sin\phi d\phi=0.$ (37)
We may now present the following physical argument: In the frame $K_{0}$, let the shell be displaced by an amount $\Delta y$ in the $y$ direction. The mechanical work spent is thus $\Delta W=F_{y}\cdot\Delta y=0$. But it is impossible that such a process where no work has been done, can lead to an electric current, producing energy. According to this, we must conclude that the process envisaged in [1, 2] cannot work. The same can be said for a stationary situation, in which the shell moves with constant velocity $v$ for an arbitrary time.
Based on our analysis, one can conclude that the process of electric power generation from Earth’s rotation cannot start at all under the given circumstances, contrary to the conclusion made in [1, 2].
It seems to us, however, that if one places the dielectric shell in another direction, which has an angle with respect to the Earth’s magnetic field, one might arrive at a different conclusion, namely to produce electric current. This case is currently under study.
Acknowledgements
We are grateful to Anca Tureanu for illuminating discussions and reading the manuscript.